(a) use the Pythagorean theorem to determine the length of the unknown side of the triangle, (b) determine the perimeter of the triangle, and (c) determine the area of the triangle.
using a right triangle, the left side is 15 in., the right side is 12 in. and the bottom is a

Answer by rfer(16322)

c^2=144+225
c^2=369
c=19.2 in
P=46.2 in